a) find and classify the critical points of the functions f(x) = 2x^3 + 3x^2 - 12 x +1 into maximum, minimum and inflection points as appreciate.

(b) The sum of two positive numbers is S. find the maximum value of their product.


1
Expert's answer
2021-07-21T17:20:38-0400

a)


"f(x)=2x^3 + 3x^2 - 12 x +1"

Domain: "(-\infin, \infin)"


"f'(x)=(2x^3 + 3x^2 - 12 x +1)'=6x^2+6x-12"

Find the critical number(s)


"f'(x)=0=>6x^2+6x-12=0"

"x^2+x-2=0"

"(x+2)(x-1)=0"

"x_1=-2, x_2=1"

Critical numbers: "-2, 1."


If "x<-2, f'(x)>0, f(x)" increases.

If "-2<x<1, f'(x)<0, f(x)" decreases.

If "x>1, f'(x)>0, f(x)" increases.


"f(-2)=2(-2)^3+3(-2)^2-12(-2)+1=21"

"f(1)=2(1)^3+3(1)^2-12(1)+1=-6"

The function "f" has a local maximum with value of "21" at "x=-2."


The function "f" has a local minimum with value of "-6" at "x=1."




"f''(x)=(6x^2+6x-12)'=12x-6"

"f''(x)=0=>12x-6=0=>x=\dfrac{1}{2}"

"f(\dfrac{1}{2})=2(\dfrac{1}{2})^3+3(\dfrac{1}{2})^2-12(\dfrac{1}{2})+1=-4"

If "x<\dfrac{1}{2}, f''(x)<0, f(x)" is concave down.

If "x>\dfrac{1}{2}, f''(x)>0, f(x)" is concave up.

Point "(\dfrac{1}{2}, -4)" is an inflection point.


Point "(-2, 21)" is a local maximum.

Point "(1, -6)" is a local minimum.

Point "(\dfrac{1}{2}, -4)" is an inflection point.



(b)

Let "x=" the first number, let "y=" the second number.

Given "x+y=S, x>0, y>0."


Then "y=S-x, 0<x<S,"


The product of these numbers is "P=xy."

Substitute


"P=P(x)=x(S-x), 0<x<S"

Find the first derivative with respect to "x"


"P'(x)=(x(S-x))'=S-x-x=S-2x"

Find the critical number(s)


"P'(x)=0=>S-2x=0=>x=\dfrac{1}{2}S"

If "0<x<\dfrac{1}{2}S, P'(x)>0, P(x)" increases.


If "\dfrac{1}{2}S<x<S, P'(x)<0, P(x)"decreases.

The function "P" has a local maximum at "x=\dfrac{1}{2}S."

Since the function "P" has the only extremum on "(0, S)," then the function "P" has the absolute maximum on "(0, S)" at "x=\dfrac{1}{2}S."


"y=S-\dfrac{1}{2}S=\dfrac{1}{2}S"

"P=\dfrac{1}{2}S(\dfrac{1}{2}S)=\dfrac{1}{4}S^2."

The maximum value of the product is "\dfrac{1}{4}S^2."